OpenStudy (anonymous):

Solve. –317 – 221 – (–354)

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

@shelbydancer What is 317 + 221 =

Directrix (directrix):

Use the Google calculator if you want.

OpenStudy (anonymous):

538

Directrix (directrix):

ok, next ...

OpenStudy (anonymous):

is ................................

Directrix (directrix):

–317 – 221 – (–354) = -538 + 354 = So, do this: 538 - 354 = ?

Directrix (directrix):

Google says: https://www.google.com/?gws_rd=ssl#q=+538+-+354+%3D

OpenStudy (anonymous):

184

Directrix (directrix):

And, then a negative sign goes out front: - 184 for the answer And, also, the Google calculator will crank out the entire thing: https://www.google.com/?gws_rd=ssl#q=%E2%80%93317+%E2%80%93+221+%E2%80%93+(%E2%80%93354)%3D